Whitney Williamson has $30,000 she would like to invest. She has a choice of two interest paying ponds, one offering 6% annual interest and the other paying 7.5% annual interest. She would like to earn $2,100 in interest in one year. If she earns more than $2,100, she would need to pay a higher income tax. How much money should she invest in each bond?
Answer by Theo(2978)

you have 2 equations that need to be solved simultaneously.
-----
x + y = 30,000
.075x + .06y = 2100
-----
multiply first equation by .06 to get:
.06x + .06y = 1800
subtract it from second equation of:
.075x + .06y = 2100 to get:
.015x = 300
divide both sides by .015 to get:
x = 20,000
this makes:
y = 10,000
-----
.075 * 20,000 + .06 * 10,000 = 1500 + 600 = 2100
-----
answer is:
she should invest $20,000 in the bond earning .075 interest rate per year and she should invest $10,000 in the bond earning .06 interest rate per year.
note:
.075 interest rate = 7.5%
.06 interest rate = 6%
-----
She will earn $2,100 in interest.
-----